Here is an example using 3 CODIS STR results. DNA evidence was found at a crime scene. Due to the poor quality of the DNA sample, only 3 of the 13 CODIS STRs could be determined. The table below lists the analytical results for the three STRs in the second and third columns (Allele A and Allele B).Here is an example using 3 CODIS STR results. DNA evidence was found at a crime scene. Due to the poor quality of the DNA samBased on the probabilities (i.e. frequencies) for the various alleles for all 3 STRS, which allele is the rarest in the sourcAnd the most common allele(s)? Select one: O a. Allele 11 at D3S1358, Allele 4 at TH01, and Allele 10 at D21S11 O b. Allele 4To calculate the RMP, we must first calculate the combined probabilities, or the genotype frequency, at each locus. Recall thTo calculate the RMP, we must first calculate the combined probabilities, or the genotype frequency, at each locus. Recall thTo calculate the RMP, we must first calculate the combined probabilities, or the genotype frequency, at each locus. Recall thFinally, we want to know what the probability is for a random individual to have these 3 genotypes simultaneously for the 3 S
Answer #1

1. The correct answer is option A, TH01.

As it has two identical alleles i.e. 4.

2. The correct answer is option A, allele 7 in D21S11.

As the probability of this particular allele is lowest, p=0.1.

3. The correct answer is option A, allele 11 at D3S1358, allele 4 at TH01, allele 10 at D21S11

As all these alleles have the same probabillity, p=0.5

4. Genotype frequency for D3S1358 = 2pq= 2(0.5x0.2)= 0.2.

5. Genotype frequency for TH01= p^2 = 0.5x0.5 = 0.25.

6. Genotype frequency for D21S11= 2pq = 2(0.1x0.5)= 0.1.

7. Probability of a random individual to have these 3 genotypes for STR loci = (0.2x0.25x0.1)= 0.005